OpenStudy (anonymous):

Solve the equation: 1/2 + 1/3 x = 2/3 x=1/2 x=1/4 x=1 x=2/3

OpenStudy (anonymous):

1/2 + 1/3 x = 2/3 1/3 x = 2/3 - 1/2 1/3x = 1/6 x = 1/2

OpenStudy (anonymous):

TaDa ^

OpenStudy (anonymous):

thank you